Systems Of Nonlinear Equations

-Solve the following equations.


1. { x+2y=0

{ x²+y²=5


2. { x+y=3

{ x²+y²=2


1
Expert's answer
2022-05-12T15:16:19-0400

1.


"\begin{cases}
 x=-2y\\
 (-2y)^2+y^2=5
\end{cases}"

"\begin{cases}
 x=-2y\\
 5y^2=5
\end{cases}"

"\begin{cases}
 x=-2y\\
 y^2=1
\end{cases}"

"\begin{cases}
 x=2\\
 y=-1
\end{cases}\text{ or }\begin{cases}
 x=-2\\
 y=1
\end{cases}"

"(-2, 1), (2, -1)"


2.


"\begin{cases}
 x=3-y\\
 (3-y)^2+y^2=2
\end{cases}"

"\begin{cases}
 x=3-y\\
 9-6y+y^2+y^2=2
\end{cases}"

"\begin{cases}
 x=3-y\\
2y^2-6y+7=0
\end{cases}"

"D=(-6)^2-4(2)(7)=-20<0"

The system has no real solution.